OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -1, 2, 5, 8, ... an = -1 + 3(n) an = -1 + 3(n - 1) an = -1 + 3 an = -1 + 3(n + 1)

OpenStudy (anonymous):

@sauravshakya @phi @Hero

OpenStudy (shubhamsrg):

plug in n=1 and see which eqn gives you ans -1 as -1 is your first term ;) that was the layman method..

OpenStudy (shubhamsrg):

otherwise,,you need to read a book on AP i guess..the first page or second page of that shall provide you the ans..

OpenStudy (anonymous):

that doesnt help me...

OpenStudy (shubhamsrg):

i see..

OpenStudy (shubhamsrg):

:P

OpenStudy (anonymous):

can you show me how to solve this?

OpenStudy (shubhamsrg):

what the nth term of an AP ? you must be knowing what an AP is right ?

OpenStudy (phi):

Before even looking at the choices, I would look at the progression and find the difference between each number.

OpenStudy (anonymous):

3

OpenStudy (phi):

with -1, 2, 5, 8, ... I would notice that 2 - -1= 3 5-2=3 8-5=3 so I expect the numbers to go up by 3 (seems obvious?)

OpenStudy (anonymous):

yes

OpenStudy (phi):

now I would look at the choices. the first one says an = -1 + 3(n) that 3 in 3n looks good. But I notice all the choices have a 3 in them... How do we pick the right choice? use shub's idea if n=1 (that means the first term, term number 1) we replace the n with 1 a1 (that just means the value of the first term) is -1 + 3*1= -1+3=2 2 does not match -1 (which is the first term in -1, 2, 5, 8, ...

OpenStudy (phi):

can you check the 2nd choice an = -1 + 3(n - 1)

OpenStudy (anonymous):

-1+3*1=-1+3=2???

OpenStudy (phi):

that is the first choice an= -1+3n check an = -1 + 3(n - 1) everywhere you see n, put in 1. then do the arithmetic (but remember an becomes a1, and that is just a name)

OpenStudy (anonymous):

a1=-1+3(1-1)

OpenStudy (phi):

yes, exactly. now what is (1-1)?

OpenStudy (anonymous):

0

OpenStudy (phi):

yes, so what is a1=-1+3(1-1)

OpenStudy (anonymous):

2

OpenStudy (phi):

?

OpenStudy (anonymous):

oh wait, 0

OpenStudy (phi):

that way you simplify a1=-1+3(1-1) is to do the subtraction in the parens (1-1). replace the (1-) with the answer 0 a1= -1 +3*0 now do 3*0 add to -1

OpenStudy (phi):

* replace the (1-1) with the answer 0

OpenStudy (anonymous):

3*0+ -1?

OpenStudy (phi):

three times zero plus -1

OpenStudy (anonymous):

-1

OpenStudy (phi):

so the 2nd choice tells you a1= -1 this formula gives -1 as the 1st term. That matches the first term of -1, 2, 5, 8, ... so this might be the answer. Check the formula with n=2 to see if you get a 2

OpenStudy (phi):

can yo do that?

OpenStudy (anonymous):

a2 = -1 + 3(2 - 1)?

OpenStudy (phi):

yes, keep going.

OpenStudy (anonymous):

2!

OpenStudy (phi):

Yes, a2= -1 +3*(2-1) 2-1 is 1 so a2= -1+3*1 a2= -1+3 a2=2 so choice 2 looks like it's the answer. you could check that choices 3 and 4 do not work, but that is not necessary, except if you want to practice.

OpenStudy (anonymous):

thank you! :)